### v2

# CaliDrop: KV Cache Compression with Calibration

Abstract
Large Language Models (LLMs) require substantial computational resources during generation. While the Key-Value (KV) cache significantly accelerates this process by storing attention intermediates, its memory footprint grows linearly with sequence length, batch size, and model size, creating a bottleneck in long-context scenarios. Various KV cache compression techniques, including token eviction, quantization, and low-rank projection, have been proposed to mitigate this bottleneck, often complementing each other. This paper focuses on enhancing token eviction strategies. Token eviction leverages the observation that the attention patterns are often sparse, allowing for the removal of less critical KV entries to save memory. However, this reduction usually comes at the cost of notable accuracy degradation, particularly under high compression ratios. To address this issue, we propose CaliDrop, a novel strategy that enhances token eviction through calibration. Our preliminary experiments show that queries at nearby positions exhibit high similarity. Building on this observation, CaliDrop performs speculative calibration on the discarded tokens to mitigate the accuracy loss caused by token eviction. Extensive experiments demonstrate that CaliDrop significantly improves the accuracy of existing token eviction methods.
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across various domains (gpt4, ; llama2, ; llama3, ; mistral, ; deepseek-v3, ). However, their auto-regressive generation process often results in slow inference speeds. KV cache reduces the computational complexity of self-attention from $O(n_{2})$ to $O(n)$ by storing intermediate key-value pairs from previous attention operations. This allows the model to avoid recalculating these values in future attention computations, thus speeding up inference. However, this efficiency comes at the cost of additional memory usage, as the KV cache must store the key-value pairs for each token. The memory overhead of the KV cache grows linearly with the sequence length, which becomes particularly severe in long-context scenarios (Streamingllm, ), posing substantial deployment challenges and creating the need for effective KV cache compression solutions.
Existing KV cache compression methods include quantization (GEAR, ; KIVI, ), layer-wise sharing (layer1, ; layer2, ), prefix sharing (batch1, ; batch2, ), head-wise sharing (MQA, ; GQA, ), token eviction (snapkv, ; H2O, ), and low-rank projection (lora1, ; lora2, ). Our work focuses on token eviction, which reduces memory consumption by removing non-critical tokens from the KV cache. Current token eviction methods typically leverage the observation that a small subset of tokens contributes to the majority of attention scores. For example, StreamingLLM (Streamingllm, ) argues that the tokens at the beginning and those closest to the current token have higher attention scores. H2O (H2O, ) estimates the importance of each token by using the accumulated attention score. SnapKV (snapkv, ) estimates token importance using an observing window and the pooled accumulated attention score. However, these approaches face two fundamental limitations in high compression ratio scenarios: (1) Permanently discarding tokens can be detrimental to accuracy if those tokens later become crucial, and (2) the coarse-grained removal process often leads to significant accuracy degradation (MiKV, ).
To address these limitations, we propose CaliDrop, a novel strategy that enhances token eviction methods through calibration. Our key insights are based on two empirical observations: (1) Queries at nearby positions exhibit high similarity, and (2) Historical attention outputs can be used to predict future attention outputs. Building on these insights, CaliDrop compensates for evicted tokens by precomputing attention outputs for queries at nearby positions, alleviating memory pressure while maintaining model accuracy. Extensive experiments demonstrate that CaliDrop significantly enhances the accuracy of existing token eviction methods.
In summary, our contributions are as follows:
  • We analyze the similarity between queries of the model and find that queries at nearby positions typically exhibit higher cosine similarity.
  • We propose CaliDrop, a general strategy that can be integrated with existing token eviction methods to effectively approximate compensation for evicted tokens, thus improving accuracy under high compression ratios.
  • Experiments on different benchmarks, models, and compression ratios have demonstrated that our method significantly improves accuracy over existing token eviction methods.
2 Related work
KV Cache Compression.  KV cache compression is essential for reducing memory usage and improving efficiency in Large Language Models, allowing faster inference and deployment on resource-constrained scenarios. There are different methods for KV cache compression. Low-rank approximation (GEAR, ; palu, ) compresses KV cache by leveraging their low-rank structure. Group Query Attention (GQA) (GQA, ) and Multi-Query Attention (MQA)  (MQA, ) share KV cache within different heads to reduce redundancy. Multi-Head Latent Attention (MLA) (deepseek, ) down-projects KV cache to low-rank spaces for memory optimization. Token eviction (H2O, ; Streamingllm, ; snapkv, ) discard some unimportant tokens during generation, in order to reduce the size of the KV cache while preserving higher accuracy as much as possible. KV cache quantization (flexgen, ; KIVI, ; KV_QUANT, ; zipcache, ; CQ, ; QAQ, ; MiKV, ) stores unimportant tokens in lower precision and retains important ones in full precision. Inter-layer KV sharing (YOCO, ; layer1, ; layer2, ) reduces memory usage by allowing different layers to share the same KV cache, significantly lowering redundancy across layers. Prefix sharing (batch1, ; batch2, ) shares common prefixes among different sequences to reduce redundancy. KV merging (kvmerger, ; cam, ) uses reparameterization and interpolation techniques to reduce redundancy while preserving semantic integrity. These approaches are often orthogonal and can typically be combined to achieve superior compression results and efficiency gains.
Token Eviction.  To reduce memory usage, eviction-based strategies keep a fixed KV cache size to store critical KV pairs and discard unnecessary pairs. Most methods evaluate token importance based on attention scores. As demonstrated in (scissorhands, ; H2O, ; snapkv, ), tokens with higher accumulated attention scores are considered more important. Alternative strategies employ factors such as initial tokens (Streamingllm, ), special tokens (FastGen, ; SepLLM, ), or L2 norms (L2Norm, ). Recent studies focus on optimizing the allocation of the KV cache memory budget. Some focus on cross-layer strategies, where PyramidKV (pyramidkv, ) and PyramidInfer (pyramidinfer, ) utilize a pyramid-shaped memory allocation while selecting tokens with high attention scores in each layer. CakeKV (cake, ) dynamically analyzes the attention patterns of each layer during the prefill stage to adaptively allocate KV cache sizes. DynamicKV (dynamickv, ) computes the average attention scores of recent and historical tokens, allocating budgets proportionally based on the density of critical tokens in each layer. Other works focus on head-level allocation strategies. AdaKV (ada, ) optimizes L1 loss between original and pruned multi-head attention outputs for head-wise budget assignment. LeanKV (leankv, ) optimizes KV cache management through heterogeneous quantization, dynamic sparsity, and unified compression. HeadKV (headkv, ) evaluates the retrieval and reasoning performance to optimize the KV cache allocation. In order to minimize redundancy and mitigate information loss due to the token eviction strategy, researchers have proposed several innovative methods. For example, CaM (cam, ) and KVMerger (kvmerger, ) merge less important tokens into a single token based on similarity in attention patterns. Additionally, RecycledAttention (recycledattention, ) recalculates previously deemed unimportant tokens at regular intervals to select the top-k important tokens.
3 Method
3.1 Preliminary Experiments
While prior studies indicate historical attention scores can approximate future patterns (H2O, ; snapkv, ), the predictive power of historical queries and the utility of historical attention outputs for future approximations remain less explored. To establish a foundation for our proposed method, we conduct preliminary experiments addressing two key questions: (1) Can historical queries approximate future queries? (2) Can historical attention outputs help approximate future attention outputs?
3.1.1 Q1: Can Historical Queries Approximate Future Queries?
To investigate whether historical queries can approximate future queries, we conduct experiments using LLaMA-3-8b-Instruct (llama3, ) on LongBench (longbench, ). We retain all queries of Layer 10, Head 16 during inference and calculate the cosine similarity between each of them. Figure 1 (top-right) presents a heatmap of the cosine similarity of queries in a single sample (we plot the results of token positions from 50-150 for simplicity). While we show only one figure for clarity, similar trends are observed across other layers, heads, tokens and samples. The heatmap indicates that queries at nearby positions exhibit high cosine similarity, suggesting that historical queries indeed provide a reliable approximation for future queries.
3.1.2 Attention Decomposition Theorem
Before answering Q2, we propose a theoretical foundation for splitting attention computation. Let $Q$ be the Query of the current tokens, and let $K$, $V$ represent the Key and Value of a set of $n$ tokens $S=\{t_{1},\dots,t_{n}\}$ in the KV cache. Consider a partition of the tokens into disjoint subsets $S_{i}$ and $S_{j}$ such that $S_{i}\cup S_{j}=S$. Let $K_{S_{i}},V_{S_{i}}$ and $K_{S_{j}},V_{S_{j}}$ be the corresponding subset of $K$ and $V$. The attention computation over the full set $S$ can be decomposed as follows:
$\text{Att}(Q,K,V)=\alpha_{i}\cdot\text{Att}(Q,K_{S_{i}},V_{S_{i}})+\alpha_{j}\cdot\text{Att}(Q,K_{S_{j}},V_{S_{j}})$ (1)
Where the weights $\alpha_{i}$ and $\alpha_{j}$ represent the normalized exponential sum of the attention weighs from each subset:
$\alpha_{i}=\frac{\sum_{t\in S_{i}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}},\quad\alpha_{j}=\frac{\sum_{t\in S_{j}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}$ (2)
Note that $\alpha_{i}+\alpha_{j}=1$. This decomposition allows for the independent attention computation of the subsets and is crucial for our hybrid attention computation strategy.
Proof.
$\text{Att}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V=\frac{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}V_{t}}{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}$ (3)
$=\frac{\sum_{t\in S_{i}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}V_{t}+\sum_{t\in S_{j}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}V_{t}}{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}$
$=\frac{\sum_{t\in S_{i}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}\cdot\frac{\sum_{t\in S_{i}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}V_{t}}{\sum_{t\in S_{i}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}+\frac{\sum_{t\in S_{j}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}{\sum_{t\in S}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}\cdot\frac{\sum_{t\in S_{j}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}V_{t}}{\sum_{t\in S_{j}}e^{\frac{QK_{t}^{\top}}{\sqrt{d_{k}}}}}$
$=\alpha_{i}\cdot\text{Att}(Q,K_{S_{i}},V_{S_{i}})+\alpha_{j}\cdot\text{Att}(Q,K_{S_{j}},V_{S_{j}})$
3.1.3 Q2: Can Historical Attention Outputs Help Approximate Future Attention Outputs?
In this part, we investigate whether historical attention outputs can be used to approximate future attention outputs. Directly multiplying the historical query with the KV cache is not feasible, as it would discard the information from the current query entirely. Therefore, we propose a method that splits the KV cache into two parts (important & unimportant) inspired by the token eviction methods. We compute the current query with the important KV cache and the historical query with the unimportant KV cache. We integrate the two results according to Equation 1 as the final attention output to approximate the true current attention output. We refer to this integration process as calibration, which is the core of our method.
To preliminarily validate the effectiveness of this method, we conduct experiments using LLaMA-3-8b-Instruct (llama3, ) on LongBench (longbench, ). We use the token eviction strategy from SnapKV (snapkv, ) to select 128 important tokens and other tokens are considered unimportant. We set the current token to be the first token after the prefill phase. We set the historical token to be one of the 0-9th tokens preceding the current token (randomly selected). We compare the L1 loss of the approximated attention output and the true current attention output before and after calibration across different layers.
Theoretically, if the historical query is completely the same as the current query, the L1 loss should be 0. When the historical and current queries are highly similar, we expect the calibration to provide a positive effect. Conversely, if the queries are not similar, there will be a side effect. Figure 1 (bottom-right) shows the L1 loss before and after calibration across different layers. The results demonstrate that adding calibration significantly reduces the L1 loss. Similar trends are observed across other tokens and samples, although additional figures are omitted for brevity. From these findings, we conclude that historical attention outputs can help approximate future attention outputs.
Refer to captionRefer to caption
Figure 1: Illustration of the proposed method and preliminary results. Left: Overview of the method in decoding. We first compute the cosine similarity between the current query and historical query, and then decide whether to perform recomputation, apply calibration, or take no action based on the similarity. Top-right: the heatmap of cosine similarity of the queries in an example. Bottom-right: L1 loss before and after calibration.
3.2 CaliDrop
Algorithm 1 CaliDrop Prefill
0: Input sequence $X$, eviction strategy $f$
1: Compute all queries $Q$, keys $K$, values $V$ for input $X$
2: Compute attention outputs for current query: $\text{Output}=\text{Att}(Q,K,V)$
3: Select important tokens using eviction strategy: $KV_{\text{compress}}\leftarrow f(Q,K,V)$
4: Offload evicted KV pairs: $KV_{\text{evict}}\leftarrow KV_{\text{full}}\setminus KV_{\text{compress}}$
5: Compute calibration using the last query:
[…]
9:return  Output, $KV_{\text{compress}}$, $KV_{\text{evict}}$, $C_{weight}$, $C_{out}$, $C_{q}$
Algorithm 2 CaliDrop Decode
0: Current query $Q_{t}$, $KV_{\text{compress}}$, $KV_{\text{evict}}$, $C_{weight}$, $C_{out}$, $C_{q}$, thresholds $\theta_{1}<\theta_{2}$
1: Compute attention: $\text{Output},A_{weight}=\text{Att}(Q_{t},K_{\text{compress}},V_{\text{compress}})$
2: Compute the similarity of the current query and the historical query: $\rho\leftarrow\cos(Q_{t},C_{q})$
3:if $\rho<\theta_{1}$ then
4:  Reload $KV_{\text{evict}}$ and perform recomputation:
5:  $C_{weight}=\sum\exp\left(\frac{Q_{-1}*K_{\text{evict}}^{\top}}{d_{k}}\right)$
6:  $C_{out}=\text{Att}(Q_{-1},K_{\text{evict}},V_{\text{evict}})$
7:  $C_{q}=Q_{t}$
8:  Add calibration with Equation 1: $\text{Output}=\alpha_{i}\cdot\text{Output}+\alpha_{j}\cdot C_{out}$
9:  return  Output
10:else if $\rho>\theta_{2}$ then
11:  Add calibration with Equation 1: $\text{Output}=\alpha_{i}\cdot\text{Output}+\alpha_{j}\cdot C_{out}$
12:  return  Output
13:else
14:  return  Output
15:end if
After the aforementioned preliminary results, we introduce CaliDrop. CaliDrop can be applied on top of existing token eviction methods to enhance accuracy. Following SnapKV snapkv , we focus on token eviction during prefilling, although our method can be easily extended to the decoding phase.
After token eviction based on a certain strategy, we do not directly delete the evicted KV cache. Instead, we offload them. At this point, we use the current query to perform attention computation with the evicted KV cache. After computation, we retain the current query, the softmax denominator, and the attention output for future calibration during subsequent attention computations. In the decoding phase, we add calibration according to Equation 1. However, the similarity between the current query and the historical query decreases as the positional difference increases. Thus, we need to ensure that the added calibration mostly provides a positive effect. To address this, we introduce two hyper-parameters: the left threshold ($\theta_{1}$) and the right threshold ($\theta_{2}$). These thresholds control whether recomputation should be performed and calibration should be applied. We calculate the cosine similarity between the current query and the retained historical query. If the cosine similarity is lower than $\theta_{1}$, we consider the current query and historical query to be sufficiently distant. In this case, we reload the evicted KV cache, recompute the attention output with the current query, and update the historical query, the softmax denominator, and the attention output. This hyper-parameter controls the frequency of recomputation, thus controlling the additional computation overhead. If the cosine similarity is higher than $\theta_{2}$, we assume the current query and the historical query are sufficiently similar, and we add calibration because we believe this calibration is likely to improve accuracy. When the cosine similarity lies between the two thresholds, no recomputation or calibration is performed. $\theta_{1}$ and $\theta_{2}$ work together to balance the accuracy of calibration and the additional computational overhead. The prefilling and decoding procedures of CaliDrop are formally described in Algorithms 1 and 2, respectively.
4 Experiments
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
MF-en
Qasper
HotpotQA
2WikiMQA
GovReport
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
18409 3619 9151 4887 8734 2113 5177 8209 6258 11141 9289 1235 4206
Mistral-7B-Instruct, KV Size = Full
FullKV 48.54 24.35 32.92 21.87 33.05 25.77 67.00 86.84 40.95 5.40 91.00 57.24 49.84 44.98
Mistral-7B-Instruct, KV Size = 64
SLM 22.76 10.31 16.59 14.09 13.41 12.21 35.67 75.93 26.47 3.66 64.60 41.27 35.41 28.64
   +CaliDrop 29.62 13.29 22.54 15.70 20.09 19.14 54.00 78.68 30.71 7.09 63.56 44.62 37.99 33.62
H2O 32.91 13.15 17.60 14.29 19.56 18.99 39.33 79.77 35.80 5.09 74.76 48.19 40.14 33.81
   +CaliDrop 37.39 17.73 22.19 16.93 23.03 22.01 56.00 83.69 37.12 5.32 77.89 49.31 42.88 37.81
SnapKV 34.46 12.87 20.11 15.52 17.33 16.37 39.00 81.14 35.14 6.89 73.55 47.05 39.80 33.79
   +CaliDrop 39.16 16.94 22.90 17.67 21.83 21.38 55.33 84.25 37.09 7.17 80.33 46.72 41.93 37.90
Mistral-7B-Instruct, KV Size = 128
SLM 22.66 8.23 18.15 14.98 14.72 14.25 41.00 79.27 34.82 3.74 40.32 48.47 39.08 29.21
   +CaliDrop 30.34 12.31 22.55 16.61 20.47 20.39 53.33 79.95 36.03 7.14 54.29 50.62 41.58 34.28
H2O 32.94 13.75 18.72 15.08 21.20 20.71 43.00 82.20 37.69 6.38 82.50 51.3 42.55 36.00
   +CaliDrop 37.74 18.60 23.47 17.62 23.78 22.37 55.67 85.54 38.88 5.95 83.92 51.50 44.86 39.22
SnapKV 37.44 14.85 21.45 15.48 20.76 20.15 47.33 84.46 37.34 5.56 85.22 51.04 42.50 37.20
   +CaliDrop 41.83 18.75 24.52 18.12 23.68 22.23 56.33 86.19 37.90 5.63 86.14 51.68 44.53 39.81
LLaMA-3-8B-Instruct, KV Size = Full
FullKV 40.56 37.54 49.81 34.93 31.04 25.58 69.67 89.85 40.50 12.94 83.67 56.58 51.01 47.98
LLaMA-3-8B-Instruct, KV Size = 64
SLM 24.73 20.83 43.86 29.93 14.59 13.50 36.67 71.49 25.80 14.67 81.00 49.67 42.81 36.12
   +CaliDrop 31.52 27.21 46.79 32.17 17.27 17.27 60.00 81.60 29.45 12.33 78.00 49.39 43.79 40.52
H2O 31.14 24.28 47.41 31.25 19.02 18.95 38.33 86.77 35.58 8.69 82.33 57.38 48.44 40.74
   +CaliDrop 35.92 28.80 49.58 32.83 20.85 20.86 62.00 88.08 36.95 11.44 82.67 55.75 47.92 44.13
SnapKV 33.33 25.05 47.56 31.83 16.79 17.20 40.67 86.06 33.81 12.22 77.67 56.27 48.07 40.50
   +CaliDrop 36.91 31.27 49.33 33.97 19.16 19.73 62.33 88.10 35.56 11.55 78.67 54.85 45.55 43.61
LLaMA-3-8B-Instruct, KV Size = 128
SLM 26.21 21.09 43.69 29.56 15.82 16.21 41.33 74.01 33.13 12.67 77.33 56.66 48.40 38.16
   +CaliDrop 30.89 26.28 47.56 31.20 17.98 19.41 61.33 84.39 34.90 12.33 77.67 56.75 46.49 42.09
H2O 36.17 25.48 48.65 31.74 20.76 20.67 40.00 87.30 36.04 12.33 84.00 58.60 51.23 42.54
   +CaliDrop 34.90 29.42 49.91 32.62 22.43 21.66 62.00 89.42 37.60 13.17 83.67 56.55 49.46 44.83
SnapKV 37.76 29.75 49.19 32.13 19.89 20.18 47.33 87.86 35.33 11.11 82.33 60.29 50.87 43.39
   +CaliDrop 37.23 30.87 50.45 33.19 21.67 21.57 65.00 89.35 36.94 12.44 83.00 56.19 49.07 45.15
Table 1: Performance comparison of CaliDrop with SnapKV, H2O, StreamingLLM (SLM) and FullKV on LongBench for LLaMA-3-8B-Instruct and Mistral-7B-Instruct. CaliDrop generally achieves improvements over previous KV cache compression methods across various KV cache sizes and LLMs. The performance strengths of CaliDrop are more evident in small KV cache sizes (i.e. KV Size = 64). Bold text represents the best performance.
4.1 Experiment Setup
4.1.1 Models and Benchmarks
We test our method on three LLMs: Mistral-7B-Instruct (mistral, ), LLaMA-3-8B-Instruct, and LLaMA-3-70B-Instruct (llama3, ). These models are chosen for their strong performance and wide utilization in natural language processing tasks.
To thoroughly evaluate the capabilities of these models with different methods, we use three benchmarks that focus on different aspects of language understanding and reasoning.
LongBench  (longbench, ): This benchmark is a comprehensive benchmark to evaluate the contextual understanding capabilities of LLMs. It includes tasks such as answering questions, summarizing, and generating code.
RULER  (hsieh2024ruler, ): RULER is a benchmark to evaluate the long-context modeling capabilities of LLMS. It involves tasks that require understanding and integrating various pieces of information, making it essential for assessing skills in complex multi-hop and aggregation tasks.
Needle-in-a-Haystack  (liu2024lost, ): This benchmark focuses on testing if models can find important details in long texts. It checks how well models can spot useful information in a lot of text, which is key for tasks like finding facts or answering questions by pulling out parts of the text.
4.1.2 Baselines
We evaluate the performance improvement of CaliDrop on top of three baselines.
StreamingLLM (SLM)  (Streamingllm, ) identifies the attention sink phenomenon and enables LLMs to process infinitely long texts by maintaining KV cache of the attention sink along with a sliding window of recent tokens.
Heavy Hitter Oracle (H2O)  (H2O, ) measures the importance of tokens using accumulated attention scores. It retains a sliding window along with the KV cache of important tokens, thereby preserving the most useful information.
SnapKV  (snapkv, ) achieves KV cache compression by selecting clustered important tokens and utilizing a local window in prefilling phase. It incorporates a clustering algorithm with a pooling layer and captures attention signals from an observation window.
FullKV caches all keys and values corresponding to each token, which is the standard approach for KV Cache in transformer-based models.
Note: For simplicity, we follow the setting of SnapKV (snapkv, ) and perform compression only in the prefilling stage. Specifically, during prefilling, token eviction is conducted separately according to the strategies of each baselines. Afterward, the decoding process proceeds with our calibration for better accuracy. Although our approach only focuses on compression during the prefilling stage, it can be easily extended to the decoding stage as well.
4.1.3 Implementation Details
We evaluate the performance of various methods under different compression ratios by retaining {64, 128, 256, 512} tokens of the prompt. For StreamingLLM, we set the number of the attention sinks to 32, and the remaining tokens are allocated as local tokens. For H2O, we evenly split the token budget into important tokens and local tokens, each comprising half of the total. For SnapKV, we set the observation window size to 32 and employ average pooling with a kernel size of 5 for importance evaluation. For CaliDrop, we set $\theta_{1}$ to 0.7 and $\theta_{2}$ to 0.85, which achieves a trade-off between accuracy improvement and computational efficiency (see Section 5 for more details and exploration experiments). All experiments are conducted on NVIDIA A100 40G GPUs.
4.2 Main Results
4.2.1 Results on LongBench
Table 1 and Appendix B show the main results of each method with different models in LongBench. We can conclude from the table that:
CaliDrop can enhance the performance of various methods: CaliDrop demonstrates significant performance improvements across a wide variety of tasks, effectively mitigating the performance degradation caused by KV cache compression. Specifically, in the Mistral-7B-Instruct and LLaMA-3-8B-Instruct models with a KV size of 64, CaliDrop consistently outperforms traditional baseline methods such as SLM, H2O, and SnapKV. This suggests that CaliDrop is highly effective in counteracting performance losses, especially under high compression ratios. Moreover, its performance gains are particularly notable in tasks that require multi-hop reasoning and multi-document summarization, highlighting its ability to maintain high performance even in complex and demanding scenarios.
Marginal benefit of CaliDrop will diminish at higher KV sizes: As the KV size approaches full capacity, the performance improvement brought by CaliDrop becomes increasingly constrained. This trend can be attributed to the fact that an already large KV cache provides ample information for the model to achieve satisfactory performance. Consequently, the additional benefits derived from calibration are diminished or even reversed. Moreover, as shown in Equation 1, when the KV size increases, $\alpha_{i}$ will also increase, while $\alpha_{j}$ will decrease, reducing the calibration gain.
4.2.2 Results on RULER
To further test our method in the long context of reasoning ability, we use the advanced long-context benchmark, RULER. In this experiment, we take Mistral-7B-Instruct and LLaMA-3-8B-Instruct as the base model and set the cache size to 64, 128, 256, and 512. As demonstrated in Appendix C, CaliDrop gains a comprehensive performance enhancement across different KV sizes. For example, with a KV size of 64, LLaMA-3-8B-Instruct using SnapKV with CaliDrop reaches 23.32% compared to the baseline (14.95%). As the KV size increases, CaliDrop continues to outperform the baselines. On Ruler, we don’t observe the gradual diminishing of the calibration gains as seen in LongBench. This is likely because the KV size of 512 was far from the required capacity for the task. Overall, the results demonstrate the effectiveness of CaliDrop.
4.2.3 Results on Needle-in-a-Haystack
Figure 2: Results of Needle-in-a-Haystack on LLaMA-3-8B-Instruct with 8k context size and 64 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Group
MF-en
Qasper
HotpotQA
2WikiMQA
GovReport
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
AVG
R=0.80 L=0.60 36.80 30.28 49.46 32.78 20.70 20.80 61.67 88.62 36.21 13.28 82.00 61.33 51.24 45.01
L=0.65 37.63 31.67 49.73 32.74 21.14 21.19 63.67 88.61 37.13 12.89 83.33 58.09 50.37 45.25
L=0.70 37.99 30.80 49.53 32.66 22.26 21.81 65.33 88.95 37.60 12.11 83.00 58.08 49.67 45.37
L=0.75 38.53 33.60 48.73 32.05 22.79 22.11 65.33 89.01 38.83 10.78 83.00 58.60 49.42 45.60
L=0.80 38.66 32.71 48.85 31.88 23.77 22.79 65.67 88.64 38.29 12.00 82.67 57.39 48.38 45.52
R=0.85 L=0.60 37.98 31.74 49.05 32.71 20.50 20.64 59.67 88.77 36.31 12.22 82.33 60.65 51.24 44.91
L=0.65 36.94 31.43 49.64 33.16 20.91 21.08 62.67 88.75 36.38 12.44 83.33 57.81 50.88 45.03
L=0.70 37.23 30.87 50.45 33.19 21.67 21.57 65.00 89.35 36.94 12.44 83.00 56.19 49.07 45.15
L=0.75 37.84 32.47 49.78 34.12 22.61 22.09 64.33 89.25 37.65 11.78 83.33 58.70 51.34 45.79
L=0.80 38.94 33.85 49.74 34.15 24.19 22.80 64.33 89.53 38.74 11.44 84.00 59.11 51.22 46.31
R=0.90 L=0.60 37.51 30.88 49.00 33.47 20.49 20.52 55.33 88.54 35.85 13.11 83.00 60.17 52.70 44.66
L=0.65 38.10 31.47 49.12 33.68 20.56 20.82 59.00 87.90 36.04 12.44 82.67 58.73 51.86 44.80
L=0.70 39.06 31.24 50.51 33.45 21.34 21.35 59.67 89.06 36.18 12.22 83.00 58.70 51.79 45.20
L=0.75 37.77 32.84 49.77 33.52 21.89 21.85 58.33 89.65 36.93 13.78 83.33 59.91 52.18 45.52
L=0.80 40.46 33.57 50.17 34.83 23.17 22.06 58.67 90.03 37.75 11.78 83.33 58.90 52.52 45.94
Table 2: The influence on accuracy of the left threshold and the right threshold. The experiments are conducted on LongBench with LLaMA-3-8B-Instruct, SnapKV, and a KV size of 128.
The results in Figure 2 and Appendix D indicate that CaliDrop achieves significant improvements based on token eviction methods across all settings in our experiment. Especially when using LLaMA-3-8B-Instruct with a KV size of 128, CaliDrop increases the accuracy of H2O from 48.5% to 83.8%. With a KV size of 256, when SnapKV is used in combination with CaliDrop, it manages to maintain almost no drop in performance in a context size of 8k. The reason for this enhancement is CaliDrop’s ability to calibrate attention outputs using historical data, which helps maintain the integrity of attention mechanisms despite the compression of the KV cache. This calibration process ensures that the model can still effectively process long contexts, thereby mitigating the performance degradation caused by token eviction.
5 Discussion
CaliDrop has two key hyperparameters: $\theta_{1}$ and $\theta_{2}$. Specifically, $\theta_{1}$ controls the frequency of recomputations, while $\theta_{2}$ determines the threshold conditions under which calibration is deemed acceptable.
These hyperparameters directly influence the model throughput and accuracy. Throughput is primarily affected by $\theta_{1}$, as it is directly tied to the frequency of recomputations. However, accuracy is influenced by both $\theta_{1}$ and $\theta_{2}$. $\theta_{1}$ affects calibration quality, with more frequent recomputations leading to better calibration. $\theta_{2}$ determines the criteria for when calibration is considered acceptable. If the conditions are too strict, the model may not fully utilize the benefits of calibration. On the other hand, if the conditions are too lenient, the model may suffer from low-quality calibration, which could reduce accuracy. Therefore, in this section, we focus on how to improve accuracy while minimizing the decrease in model throughput.
5.1 Influence on Accuracy
(a) Throughput
(b) Recomputation frequency
(c) Accuracy and throughput
Figure 3: (a) Throughput of CaliDrop and SnapKV under different batch size. (b) Recomputation frequency of CaliDrop in different layers under different $\theta_{1}$.(c) Accuracy and throughput under different calibration sizes. No cali: without calibration. Full cali: all evicted tokens are used for calibration (calibration size = $\infty$). We set $\theta_{1}=0.7$ in our experiment.
Table 2 shows the influence on accuracy of $\theta_{1}$ and $\theta_{2}$. We can conclude that, with a fixed $\theta_{2}$, higher $\theta_{1}$ lead to increased accuracy, which aligns with our hypothesis. This suggests that by increasing the frequency of recomputations, the quality of calibration improves, thereby enhancing accuracy. When $\theta_{1}$ is fixed and $\theta_{2}$ is varied, accuracy does not follow a consistent trend. This occurs because a high $\theta_{2}$ can discard some effective calibration, while a low $\theta_{2}$ may introduce potentially harmful calibration. This behavior is also consistent with our hypothesis. Based on these results, we set $\theta_{1}=0.7$ and $\theta_{2}=0.85$ for our main experiment.
5.2 Speed Analysis
We evaluate throughput across varying batch sizes using an input length of 1024, output length of 128, and a KV cache budget of 128 on an 80GB A800. As shown in Figure 3(a), SnapKV achieves notable throughput gains over FullKV due to its reduced memory footprint. CaliDrop adds only marginal computational overhead compared to SnapKV, while maintaining high throughput across practical batch sizes. Notably, its calibration-based accuracy improvement incurs minimal latency impact, showing that CaliDrop effectively balances efficiency and performance.
We also investigate the frequency of recomputation. We conduct experiments on LongBench using LLaMA-3-8B-Instruct with SnapKV and a KV size of 128. We compare the recompute frequency (the ratio of recomputations to total decoding steps) in different layers under different $\theta_{1}$ in Figure 3(b). We find that the degree of variation in query similarity between different layers is inconsistent, leading to differences in recompute frequencies across layers. Additionally, as observed earlier, a higher $\theta_{1}$ leads to more recomputations. Overall, in our main experiment ($\theta_{1}$=0.7), recomputation occurs approximately every eight steps.
5.3 Efficiency Optimization
In our main experiment, we did not apply any efficiency optimizations. However, various techniques can be used to improve efficiency. In principle, most existing KV cache optimization methods—such as quantization and token eviction—are applicable to the offloaded cache. Additional strategies like specialized CUDA kernels, page-level I/O (quest, ), and communication overlapping may also benefit CaliDrop. Since the theoretical computation overhead is low (recomputation occurs approximately every 8 steps), we leave extensive efficiency optimization to future work. As a lightweight yet effective enhancement, we introduce a hyper-parameter called calibration size, which controls the number of offloaded tokens by ranking them according to importance. This approach significantly reduces calibration overhead while preserving most of the accuracy benefits. Figure 3(c) shows the accuracy of LLaMA-3-8B-Instruct using SnapKV and CaliDrop on LongBench under different calibration sizes, along with throughput results from the experiment setup in Section 5.2. We find that a small calibration size can lead to significant performance improvements with only a modest sacrifice in throughput.
6 Conclusion
In this work, we propose CaliDrop, a strategy that enhances token eviction methods at the cost of additional computations. We leverage the similarity between queries to add precomputed calibration for future queries, and offload the evicted tokens for future recomputation. Calidrop effectively improves accuracy based on token eviction methods, allowing them to perform better in scenarios with lower compression ratios. We also explore a series of efficiency-related issues concerning CaliDrop and propose several potential optimization solutions.
Appendix A Limitations
Although CaliDrop can significantly improve the accuracy of existing token eviction methods, it still has some limitations:
  • The performance improvement brought by CaliDrop comes at the cost of throughput. It can be seen as an intermediate state between KV cache offloading and KV cache compression. It achieves a balance between accuracy and efficiency through offloading and infrequent recomputations. In Discussion, we analyze its efficiency in detail and explore several optimization directions, but leave more optimization methods for future work.
  • When the KV cache budget is very high, the token eviction method may already achieve accuracy close to that of Full KV. In such cases, CaliDrop cannot further improve accuracy because its upper bound is Full KV. Like we have already discussed in the analysis of LongBench, as the KV size increases, the marginal benefits of CaliDrop diminish.
  • For simplicity, we only consider the setting of SnapKV, where compression is applied solely during the prefilling phase. Although CaliDrop can be easily extended to the decoding phase, addressing efficiency issues in scenarios like long-generation still requires more refined design of the method.
Appendix B More results on LongBench.
Table 3 shows the additional results of LongBench.
Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
MF-en
Qasper
HotpotQA
2WikiMQA
GovReport
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
18409 3619 9151 4887 8734 2113 5177 8209 6258 11141 9289 1235 4206
Mistral-7B-Instruct, KV Size = Full
FullKV 48.54 24.35 32.92 21.87 33.05 25.77 67.00 86.84 40.95 5.40 91.00 57.24 49.84 44.98
Mistral-7B-Instruct, KV Size = 256
SLM 24.00 9.30 17.54 14.58 17.21 16.87 48.33 79.98 38.11 4.47 42.48 52.50 40.97 31.26
   +CaliDrop 29.26 13.04 22.63 16.25 21.91 21.60 54.67 80.06 38.29 6.24 59.63 53.58 43.33 35.42
H2O 34.18 14.97 19.83 15.66 22.54 21.93 48.00 85.54 39.01 5.99 83.40 54.67 45.42 37.78
   +CaliDrop 39.53 18.27 24.62 17.54 24.74 23.17 58.33 85.35 39.31 6.06 85.70 53.05 47.17 40.22
SnapKV 42.87 17.33 24.65 16.65 22.78 22.23 58.00 86.77 38.67 6.40 90.57 54.91 45.69 40.58
   +CaliDrop 43.97 22.32 26.46 18.66 24.94 23.11 57.33 86.78 39.23 6.00 90.08 54.56 46.97 41.57
Mistral-7B-Instruct, KV Size = 512
SLM 24.97 10.47 18.02 14.76 21.20 20.03 53.67 82.12 39.07 5.47 42.49 55.37 43.13 33.14
   +CaliDrop 31.47 13.97 24.15 16.18 23.99 22.49 55.33 82.70 39.38 6.67 59.86 55.11 45.59 36.68
H2O 36.32 16.50 21.03 16.29 24.19 23.24 53.67 86.58 39.52 6.19 86.37 55.7 47.09 39.44
   +CaliDrop 41.43 20.29 26.27 17.73 25.90 24.13 56.67 86.56 39.83 6.63 86.88 55.42 46.83 41.12
SnapKV 45.24 20.09 26.64 18.27 25.18 23.25 62.67 86.51 39.45 5.73 92.22 57.09 47.85 42.32
   +CaliDrop 45.03 22.27 29.66 20.73 26.60 23.81 59.00 86.86 39.21 5.77 92.33 55.56 47.84 42.67
LLaMA-3-8B-Instruct, KV Size = Full
FullKV 40.56 37.54 49.81 34.93 31.04 25.58 69.67 89.85 40.50 12.94 83.67 56.58 51.01 47.98
LLaMA-3-8B-Instruct, KV Size = 256
SLM 26.78 22.31 42.52 29.17 18.51 19.29 50.67 80.23 36.52 15.00 78.33 58.52 51.56 40.72
   +CaliDrop 33.11 25.70 47.92 31.40 20.06 21.31 63.33 84.59 37.44 12.69 78.67 57.55 48.47 43.25
H2O 35.78 28.26 48.71 31.51 21.78 21.44 45.33 89.45 37.92 12.11 83.67 60.02 51.56 43.66
   +CaliDrop 36.04 30.67 49.36 33.27 23.03 22.60 63.00 90.06 38.69 12.94 83.67 58.03 52.24 45.66
SnapKV 38.45 30.77 49.70 33.80 22.15 21.71 57.00 89.28 36.79 12.11 84.00 60.22 53.11 45.31
   +CaliDrop 39.11 32.72 50.08 33.63 23.49 22.54 66.00 89.99 38.04 13.33 83.67 56.57 51.36 46.19
LLaMA-3-8B-Instruct, KV Size = 512
SLM 27.12 22.83 43.69 29.94 21.90 21.94 56.67 82.81 37.81 14.31 76.33 60.62 53.25 42.25
   +CaliDrop 33.05 27.57 47.96 32.54 22.70 22.85 65.67 86.73 38.12 12.48 78.67 58.26 49.00 44.28
H2O 37.64 30.66 50.20 32.93 23.54 22.91 54.00 89.70 38.89 10.61 83.67 59.10 52.59 45.11
   +CaliDrop 38.17 32.68 49.77 33.43 24.63 23.42 64.67 90.06 39.08 11.50 83.67 59.77 51.93 46.37
SnapKV 38.34 33.65 50.66 34.62 24.48 23.06 65.67 90.03 38.47 12.56 83.67 59.90 52.02 46.70
   +CaliDrop 39.30 34.99 50.23 34.86 25.21 23.47 67.00 90.14 38.58 13.11 84.00 59.28 52.42 47.12
LLaMA-3-70B-Instruct, KV Size = Full
FullKV 47.44 44.15 62.12 54.97 31.52 25.21 71.33 91.24 43.56 22.33 89.33 61.71 63.30 54.48
LLaMA-3-70B-Instruct, KV Size = 64
SLM 30.69 24.51 53.82 50.26 15.28 14.58 38.00 86.66 28.53 22.67 89.00 53.77 46.60 42.64
   +CaliDrop 39.31 31.52 54.58 50.95 17.85 18.45 59.67 85.03 30.83 22.67 89.00 50.82 50.45 46.24
H2O 42.98 30.01 58.08 52.90 20.45 20.87 42.00 89.74 38.49 22.44 88.67 56.72 54.55 47.53
   +CaliDrop 44.06 35.00 58.12 52.11 21.51 21.53 62.00 90.55 39.65 22.44 88.67 58.37 55.91 49.99
SnapKV 42.35 31.78 58.51 53.62 18.44 19.15 44.00 89.41 37.12 22.67 88.33 57.42 54.10 47.45
   +CaliDrop 43.23 34.60 58.22 51.73 20.03 20.52 64.00 91.18 38.91 23.00 88.67 56.90 54.59 49.66
Table 3: Performance comparison of CaliDrop with SnapKV, H2O, StreamingLLM (SLM) and FullKV on Long- Bench for LLaMA-3-8B-Instruct, Mistral-7B-Instruct and LLaMA-3-70B-Instruct. CaliDrop generally achieves improvements over previous KV cache compression methods across various KV cache sizes and LLMs. Bold text represents the best performance.
Appendix C More results on Ruler
Table 4 shows the results of Ruler.
Method Single NIAH Multi-key NIAH MQ-NIAH MV-NIAH CWE FWE VT Avg.
S-NIAH-1
S-NIAH-2
S-NIAH-3
MK-NIAH-1
MK-NIAH-2
MK-NIAH-3
LLaMA-3-8B-Instruct, KV Size = Full
FullKV 100.00 98.20 97.00 99.20 91.60 95.80 99.75 97.45 97.82 82.27 98.28 96.12
LLaMA-3-8B-Instruct, KV Size = 64
SLM 0 0 0 0 0 0 0 0 0.12 28.80 0 2.63
   +CaliDrop 1.60 1.20 0 2.20 0.40 0 2.20 1.80 0.58 31.33 4.00 4.12
H2O 20.40 20.80 0 5.40 0.40 0 0.20 0.15 4.62 0.20 2.96 5.01
   +CaliDrop 76.60 40.20 0 17.00 7.80 0 15.30 14.50 12.16 3.60 11.24 18.04
SnapKV 58.40 67.20 0 21.40 13.80 0 0.75 0.35 0.16 0.20 2.24 14.95
   +CaliDrop 77.60 68.20 0 40.20 32.00 0 12.85 12.70 1.18 2.07 9.68 23.32
LLaMA-3-8B-Instruct, KV Size = 128
SLM 0.40 1.20 2.40 3.00 0.40 0 2.45 2.45 6.84 52.20 0.64 6.54
   +CaliDrop 2.40 2.80 2.40 3.00 0.60 0 5.00 4.00 13.72 51.27 5.96 8.29
H2O 41.00 43.80 2.40 15.00 2.40 0 2.30 0.30 18.00 15.13 7.96 13.48
   +CaliDrop 73.60 61.00 2.40 26.20 15.60 0 30.85 22.60 36.94 23.87 16.80 28.17
SnapKV 98.80 87.00 0 60.80 33.80 0 17.15 7.45 4.60 31.33 13.20 32.19
   +CaliDrop 99.00 87.20 0 65.80 45.40 0 49.40 40.50 13.94 39.07 24.28 42.24
LLaMA-3-8B-Instruct, KV Size = 256
SLM 1.40 1.20 2.40 3.00 1.40 1.20 2.45 2.45 17.42 75.60 3.12 10.15
   +CaliDrop 2.80 2.80 2.40 4.40 1.60 1.20 5.40 4.05 25.48 70.53 8.48 11.74
H2O 66.60 56.60 2.40 25.00 15.20 0 8.90 1.30 31.64 49.93 20.68 25.30
   +CaliDrop 83.20 68.00 2.40 40.60 20.80 0 42.70 29.80 58.18 55.20 38.76 39.97
SnapKV 100.00 92.60 0 90.20 26.00 0 76.40 36.50 13.40 47.67 91.48 52.20
   +CaliDrop 100.00 93.20 0 89.80 36.60 0 87.15 73.90 40.90 57.33 94.08 61.18
LLaMA-3-8B-Instruct, KV Size = 512
SLM 4.00 6.20 8.00 7.00 5.20 3.80 5.25 6.80 35.66 74.00 7.68 14.87
   +CaliDrop 6.20 7.40 8.00 8.60 5.60 3.80 8.20 8.00 41.38 75.67 12.64 16.86
H2O 88.20 65.60 2.40 44.00 25.60 1.40 24.30 3.00 46.08 64.33 69.28 39.47
   +CaliDrop 93.80 75.80 2.40 53.80 23.40 1.40 59.45 39.50 75.18 70.07 81.84 52.42
SnapKV 100.00 93.60 1.60 96.80 31.00 0.20 91.10 71.65 30.50 56.20 94.36 60.64
   +CaliDrop 100.00 94.20 1.40 97.20 37.40 0.20 95.55 86.50 67.24 67.47 96.40 67.60
Mistral-7B-Instruct, KV Size = Full
FullKV 100.00 100.00 96.40 94.60 99.00 80.60 95.75 92.85 83.72 83.93 96.48 93.03
Mistral-7B-Instruct, KV Size = 64
SLM 0 0 0 0 0 0 0 0 0.28 6.13 0 0.58
   +CaliDrop 0.20 0 0 0 0 0 0 0 1.14 8.20 2.32 1.08
H2O 0 0 0 0 0 0 0 0 3.50 0.07 5.00 0.78
   +CaliDrop 6.60 1.80 0 0.20 0.20 0 0.10 0.15 12.30 2.27 12.92 3.32
SnapKV 0.20 0.40 0 0.20 0 0 0 0 0.80 0.07 3.20 0.44
   +CaliDrop 9.20 4.40 0 2.60 1.40 0 0.10 0.15 6.38 1.20 8.60 3.09
Mistral-7B-Instruct, KV Size = 128
SLM 0.40 1.20 2.20 3.00 0.40 0 2.40 2.25 4.26 11.47 0.56 2.56
   +CaliDrop 0.40 1.20 2.20 3.00 0.40 0 2.15 2.00 26.96 12.93 2.96 4.93
H2O 2.20 2.20 0.20 3.20 0.20 0 0 0.30 14.40 13.47 11.40 4.32
   +CaliDrop 13.20 6.40 0.80 4.40 0.60 0 0.05 1.30 23.10 12.47 25.16 7.95
SnapKV 64.20 23.00 0 8.80 4.00 0 0.05 0.05 2.68 41.20 16.36 14.58
   +CaliDrop 79.40 40.00 0 17.60 8.20 0 0.60 3.15 8.26 43.47 27.72 20.76
Mistral-7B-Instruct, KV Size = 256
SLM 1.40 1.20 2.20 3.00 1.00 0.60 2.35 2.30 9.56 29.93 2.40 5.09
   +CaliDrop 1.40 1.20 2.20 3.00 1.00 0.60 2.25 1.85 34.82 32.27 4.60 7.74
H2O 29.40 5.80 2.00 4.20 0.40 0 1.45 0.50 30.00 46.87 22.72 13.03
   +CaliDrop 44.80 16.80 2.20 7.80 1.00 0 1.85 1.80 38.90 36.07 42.12 17.58
SnapKV 93.60 54.40 0 35.60 8.60 0 1.55 0.65 8.54 62.60 80.64 31.47
   +CaliDrop 95.20 59.80 0.20 43.60 16.20 0 8.50 8.00 23.22 70.27 84.24 37.20
Mistral-7B-Instruct, KV Size = 512
SLM 3.80 5.60 5.60 5.40 4.00 3.00 5.05 4.40 18.94 58.80 7.04 11.06
   +CaliDrop 3.80 5.60 5.60 5.40 4.00 3.00 5.00 4.00 39.12 53.93 8.28 12.52
H2O 74.80 21.00 2.00 9.60 2.20 0.60 1.75 0.50 37.66 65.87 49.24 24.11
   +CaliDrop 80.40 33.00 1.80 16.60 4.40 0.60 3.80 1.80 49.04 64.20 67.00 29.33
SnapKV 98.00 75.80 0.60 57.40 20.8 0 10.45 3.60 17.36 71.33 92.68 40.73
   +CaliDrop 98.00 77.20 2.80 60.40 28.40 0.60 28.50 13.20 40.78 78.67 92.60 47.38
Table 4: Performance comparison of CaliDrop with SnapKV, H2O, StreamingLLM (SLM) and FullKV on RULER for LLaMA-3-8B-Instruct. Bold text represents the best performance.
Appendix D More results on Needle-in-a-Haystack.
Figure 4, 5, 6, 7, 8, 9, 10, 11 show the full results of Needle-in-a-Haystack.
Figure 4: Results of Needle-in-a-Haystack on Mistral-7B-Instruct and LLaMA-3-8B-Instruct with 32k and 8k context size and full KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 5: Results of Needle-in-a-Haystack on Mistral-7B-Instruct with 32k context size and 64 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 6: Results of Needle-in-a-Haystack on Mistral-7B-Instruct with 32k context size and 96 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 7: Results of Needle-in-a-Haystack on Mistral-7B-Instruct with 32k context size and 128 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 8: Results of Needle-in-a-Haystack on Mistral-7B-Instruct with 32k context size and 256 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 9: Results of Needle-in-a-Haystack on LLaMA-3-8B-Instruct with 8k context size and 96 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 10: Results of Needle-in-a-Haystack on LLaMA-3-8B-Instruct with 8k context size and 128 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.
Figure 11: Results of Needle-in-a-Haystack on LLaMA-3-8B-Instruct with 8k context size and 256 KV size. The vertical axis of the figure represents the depth percentage, and the horizontal axis represents the token length.